OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x –2 0 4 f(x) 1 5 –59

OpenStudy (anonymous):

y = ax^2 + bx + c we must replace each pair of values and form equations

OpenStudy (anonymous):

1 = a(-2)^2 + b(-2) + c 5 = a(0)^2 + b(0) + c _59 = a(4)^2 + b(4) + c

OpenStudy (anonymous):

from the second equation c = 5 then 1 = 4a - 2b + 5 ---------> 4a - 2b = -4 ( 1) -59 = 16a + 4b + 5 -------> 16a + 4b = -64 (2) we solve multiply (1) by 2 8a - 4b = -8 16a + 4b = -64 We add (1) + (2) 24a = -72 ---------> a = -3 then we replace "a" in (1) 4(-3) - 2b = -4 -12 -2b = -4 -12 + 4 = 2b -8 = 2b b = -4

OpenStudy (anonymous):

Finally y = -3x^2 -4x + 5

OpenStudy (anonymous):

It was a pleasure to help you

OpenStudy (anonymous):

thanks!